Father's age is equal to the sum of the ages of his five children. After 15 years, his age will be only
half of the sum of the children's ages. How old is the father ?

Answer by ankor@dixie-net.com(22740)

Father's age is equal to the sum of the ages of his five children.
After 15 years, his age will be only half of the sum of the children's ages.
How old is the father ?
:
let x = age of the father, and the sum of the 5 kids
:
After 15 years, his age will be only half of the sum of the children's ages.
Each of the 5 kids will be 15 yrs older also, therefore
x + 15 = .5(x + 5(15))
x + 15 = .5(x + 75)
x + 15 = .5x + 37.5
x - .5x = 37.5 - 15
.5x = 22.5
x = 2(22.5)
x = 45 is Dad's present age